OpenStudy (anonymous):

Use the chain rule, in Leibniz notation, to find dy/dx y= u(u^2 +3)^3 , u= (x+3)^2 , x = -2

OpenStudy (anonymous):

\[\frac{dy}{du}=(u^2+3)^3+3u(u^2+3)^2\times 2u\] \[\frac{dy}{du}=(u^2+3)^3+6u^2(u^2+3)^2\] by the chain rule and the product rule

OpenStudy (anonymous):

yes i got this too

OpenStudy (anonymous):

you can also write this as \[(u^2+3)^2 (7 u^2+3)\] if you like

OpenStudy (anonymous):

\[u=(x+3)^2\] \[\frac{du}{dx}=2(x+3)=2x+6\]

OpenStudy (anonymous):

multiply together, and replace x by -2

OpenStudy (anonymous):

umm,that's exactly what i did...but somehow i got a really big number..

OpenStudy (anonymous):

looks like it would be really big to me

OpenStudy (anonymous):

cuz u = (x+3)^2 , when i sub it in, it will be ((x+3)^4 +3 )^2 well, the answer is only 540 or so

OpenStudy (anonymous):

it's 320

OpenStudy (anonymous):

ok if x = -2, then \[ u = (-2+3)^2=1\] so maybe not so big

OpenStudy (anonymous):

\[(u^2+3)^2 (7 u^2+3)\] \[4^2(10)=160\] then multiply by \[2(-2)96=-2\]

OpenStudy (anonymous):

sorry i mean \[2\times -2+6=2\]

OpenStudy (anonymous):

oh..so u sub in x first before multiplying?? then i think i've made a mistake when i combine both equations! ok thxx i will do it ur way then :)

OpenStudy (anonymous):

i get 320 if i am not mistaken

OpenStudy (anonymous):

yeah that is way easier